How many pounds of sunflower seed ($.70/lb) must be mixed with raisins & nuts ($1.90/lb) to form 40 pounds of a mixture to sell at $1.00/lb?
Answer by mananth(11019)

----------------- price ----------------quantity
sunflower seeds 0.7 ----------------x pounds
Raisins & nuts 1.9 ----------------40-x pounds
Mixture 1.00 ---------------- 40

0.7x+1.9(40-x)=1*40

0.7x+76-1.9x=40
0.7x-1.9x=40-76
-1.2x =-36
/ -1.2
x=30 pounds $0.70 sunflower seeds
10 pounds $1.90 Raisins & nuts
m.ananth@hotmail.ca